OpenStudy (anonymous):

PLEASE HELP Find the simplified form of the expression. Give your answer in scientific notation. (9x10^5)(6x10^-7)

OpenStudy (agent0smith):

You can change the order to make it easier\[\Large (9\times 10^5)(6\times 10^{-7}) = (9 \times 6) \times (10^5 \times 10^{-7})\]Try simplifying it more from here.

OpenStudy (anonymous):

Wow that really just confused me hahahah

OpenStudy (anonymous):

A) 5.4x10^-34 B) 1.5x10^-1 C) 5.4x10^-1 D) 1.5x10^-34

OpenStudy (agent0smith):

You can change the order of multiplication... that's all i did.

OpenStudy (agent0smith):

\[\Large (9 \times 6) \times (10^5 \times 10^{-7}) \]can you try to simplify this...?

OpenStudy (agent0smith):

@sheaalicesmith ...?